OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1

OpenStudy (anonymous):

-2x+y=1 | -2x+y=1 -2(0)+y=1 | -2x+0=1 y=1 |-2x=1 |-2x div -2 = x | 1 div -2 = -0.5 y=1 x= -0.5